OpenStudy (anonymous):

Which combination of integers can be used to generate the Pythagorean triple 7,24,25

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the question is a little unclear, but maybe it is this: all triples can be found by taking whole number \(m,n\) and computing \[m^2-n^2, 2mn, m^2+n^2\] is that what you have to find, \(m\) and \(\)?

Elsa213 (elsa213):

Welcome to Openstudy! :D

OpenStudy (misty1212):

@Elsa213 thanks dear, welcome to you as well!

Elsa213 (elsa213):

Thank you cx

OpenStudy (anonymous):

its apex so the answers would be like A.x=4,y=3 B.X=1, y+3 C.x=3,y=2 D.x=2 y=2

OpenStudy (misty1212):

what does "it is apex" mean?

OpenStudy (anonymous):

and thanks for the warm welcome and its an online class

OpenStudy (misty1212):

ok it looks like they used \(x\) and \(y\) whereas i used \(m\) and \(n\) but it is the same thing

OpenStudy (anonymous):

what would be the answer?

OpenStudy (misty1212):

you have \[x^2-y^2=7\] that is the same as \[(x-y)(x+y)=7\] there is only one way to factor \(7\) as \(1\times 7\) so \[x-y=1\\ x+y=7\] and using that you can solve for \(x\) and \(y\)

OpenStudy (anonymous):

so i basically plug in and solve

OpenStudy (misty1212):

you can solve that in your head: two numbers that add to 7 and are one apart

OpenStudy (anonymous):

basicaly A

OpenStudy (misty1212):

basically A?

OpenStudy (anonymous):

the answer?

OpenStudy (misty1212):

ooh i see A is \(x=4,y=3\) yes, that is right